(b) For the given sequence 1/n, 1/2n, 1/4n,…  Find the 16th term. Find sum of first sixteen terms where n is your arid number i.e. 19-arid-234 take n=234


1
Expert's answer
2021-08-02T14:37:45-0400

Given sequence is

"\frac{1}{n}, \frac{1}{2n}, \frac{1}{4n}..."

To find the 16th term and sum of first sixteen terms where n = 234.

We have the given sequence is

"\frac{1}{n}, \frac{1}{2n}, \frac{1}{4n}..."

Now let us consider the terms of the sequence to be named as ak

Then

"a_1= \frac{1}{n}, a_2 = \frac{1}{2n}, a_3=\frac{1}{4n}"

Next

"\frac{a_2}{a_1}= \frac{ \frac{1}{2n} }{\frac{1}{n}} \\and \\\frac{a_3}{a_2}=\frac{\frac{1}{4n}}{\frac{1}{2n}}=\frac{1}{2} \\"

hence the common ratio of the terms of the given sequence are equal so the given sequence is a geometric sequence whose first term is "a_1 = \frac{1}{n}" and common ratio r is "r=\frac{1}{2}"

As the given sequence is a geometric sequence so its kth term is given by "a_k=a_1r^{k-1}" i.e "a_k=\frac{1}{n}\Big(\frac{1}{2}\Big)^{k-1}"

For the 16th term we consider k=16 in "a_k=\frac{1}{n}\Big(\frac{1}{2}\Big)^{k-1}" then we have

"a_{16}=\frac{1}{n}\Big(\frac{1}{2}\Big)^{16-1} \\

a_{16}=\frac{1}{n}\Big(\frac{1}{2}\Big)^{15} \\

a_{16}= \frac{1}{32768n}"

Next the sum of first sixteen terms of the given sequence will be given by "S_{16}=\frac{a_1(1-r^{16})}{1-r}" as the formula for the sum of the first k terms with first term a1 and common ratio r<1 is given by "S_{k}=\frac{a_1(1-r^{k})}{1-r}"

For "a_1 = \frac{1}{n}" and "r=\frac{1}{2}" in "S_{16}=\frac{a_1(1-r^{16})}{1-r}" we get

"S_{16}=\frac{\frac{1}{n}(1-(\frac{1}{2})^{16})}{1-\frac{1}{2}} \\

S_{16}=\frac{\frac{1}{n}(\frac{65535}{65536})}{\frac{1}{2}} \\

S_{16}=\frac{65535}{32768n}"

So the 16th term and sum of first sixteen terms of the sequence "\frac{1}{n}, \frac{1}{2n}, \frac{1}{4n}…" are "\frac{1}{32768n}" and "\frac{65535}{32768n}" respectively.

As n = 234 so substituting n = 234 in the answers we get the 16th term to be "\frac{1}{7667712}" and sum of first sixteen terms to be "\frac{21845}{2555904}"

Hence the required answers for n = 234 are 16th term is "\frac{1}{7667712}" and sum of first sixteen terms is "\frac{21845}{2555904}"